the dissociation constant of a substituted benzoic acidat25c is 1.0*10^-4.The pH of a 0.01 M solution of its sodium salt is

Grade:11

1 Answers

Sunil Kumar FP
9 years ago
Ka (C6H5COOH) = 1 × 10−4
pH of 0.01 M C6H5COONa
C6H5COO- +H2O ----->C6H5CCOH +OH-
.01(1-h) .01h .01h
Kh=Kw/Ka
=.01h^2/(1-h)
10^-14/10^-4 =10^-2 h^2/(1-h)
[OH-]=10^-6
[OH−] = 0.01 h = 0.01 × 10−4 = 10−6
[H+] = 10−8
pH = 8